On January 1, 2020, Better, Inc. entered into an equipment lease with Canyon Corp. under which Better agrees to lease equipment that was manufactured by Canyon and that has an expected useful life of 4 years. The cost to manufacture the equipment was $500,000 and its normal sales price is $600,000. The lease term is 3 years and Canyon expects to recover the equipment’s normal sales price through 3 lease payments in order to earn an 8% rate of return. The residual value is expected to be $53,000. Better doesn't guaranteed the residual value. The lease payment is due at lease signing date and each of the following December 31. Better is aware of Canyon’s expected rate of return and normally amortizes the assets by the straight-line method. Canyon is reasonably confident that Better will make the lease payments and has no material uncertainties.

Round your numbers to the nearest whole numbers.

Required:      

  1. Calculate the annual lease payment.
  2. According to the FASB, how should the lease be classified by both Better (lessee) and Canyon (lessor)? Why?
  3. Calculate the present value of lease liability and lease receivable on lease signing date.
  4. Prepare lease amortization schedules up to 12/31/2020 for both better and Canyon.

  5. Prepare the journal entries to record the inception of the lease and the first lease payment on January 1, 2020 for both Better and Canyon.
  6. Prepare the appropriate journal entries for both Better and Canyon at December 31, 2020.
  7. Suppose at the end of the lease term, the actually residual value is $40,000 when the equipment is reverted back to lessor . State whether or not there will be any gain or loss occur to lessee and lessor.
Answer #1

Part 1 to 3

Year PV factor @ 8% Remarks
0    1.00000
1    0.92593 = 1 / 1.08
2    0.85734 = 0.92593 / 1.08
3    0.79383 = 0.85734 / 1.08
Total (0 to 2)    2.78326
Normal sales price $        600,000
Less: Present value of unguaranteed residual value (0.79383*53000) $          42,073
Amount recover through annual lease payments $        557,927
Divided : Total Present Value Factor (as above)            2.78326
Annual lease payments $        200,458
If you meet any criteria of below, then considered as a finance lease, otherwise considered as an operating lease.
Ownership criteria is ownership transferred to the lessee at the end of lease period? No
Specialized nature criteria is leased asset have special nature and which is no alternative use to the lessor? No
BPO criteria is it reasonably certain to exercise of bargain purchase option? No
Lease term criteria is the lease period is equal to or more than 75% of the economic life of the leased asset? (3/4 = 75%) Yes
Present value criteria is the present value of payments equal to or more than 90% of the fair value of the leased asset? (557927/600000 = 92.99%) Yes
Type of Lease Finance Lease
Finance lease is also known as capital lease.

Normal selling price is more than lessor's origional cost.

This lease is considered for sales type lease with profit.

Lease be classified by
Better (lessee) Capital lease
Canyon (lessor) Sales-type lease (with profit)
Present value of lease liability (200458*2.78326) $        557,927
Present value of lease receivable $        600,000

Part 4 to 6

XUS 2020-09(2) SEP.xlsx - Microsoft Excel ? X FILE HOME INSERT PAGE LAYOUT FORMULAS DATA REVIEW VIEW DEVELOPER Sign in R30 :XIA 5 2020-09(2) SEP.xlsx - Microsoft Excel ? X FILE HOME INSERT PAGE LAYOUT FORMULAS DATA REVIEW VIEW DEVELOPER Sign in R45

Part 7 last part

There is no gain or loss occur to lessee. Because of the lessee doesn't guaranteed the residual value.
Expected Residual value $      53,000
Actual Residual value $      40,000
loss occur to lessor $      13,000
Loss occur to the lessor = $13000